weight of 1 atom of an element is 6.644 into 10 power -23 gram .Calculate gram atom of element in 40 kg
1
Expert's answer
2012-11-06T10:01:25-0500
We must know how many atoms is in 40 kg.
N = 40000 g / 6.644 * 10^ -23 = 6020 * 10^23 atoms
gramatom is 6020*10^23 / 6.02*10^23 = 1000 g atom